OpenStudy (anonymous):

Solve. x2 + 8x + 25 = 0

OpenStudy (anonymous):

Do the Completing the Square method! :) it's pretty easy First you subtract 25 from both sides to get the numbers with variables by themselves.

OpenStudy (anonymous):

x2 + 8x = -25

OpenStudy (anonymous):

Then you divide 8 by 2 to get 4. Then you have to square 4. You get 16 and you have to add 16 to both sides to equal out the equation.

OpenStudy (anonymous):

x2 + 8x + 16 = -25 + 16 Simplify. x2 + 8x + 16 = -9

OpenStudy (anonymous):

It's asking for these answers –4 ± i –3 ± 4i –4 ± 3i –2 ± 4i

OpenStudy (anonymous):

Now factor the trinomial into a binomial. (x + 4)2 = -9 Then you have to find the square root of these. you end up with x + 4 = √-9 That's where the imaginary number part comes in.

OpenStudy (anonymous):

√-9 becomes 3i. x + 4 = 3i then you just simplify that

OpenStudy (anonymous):

so x = -4 + 3i

OpenStudy (anonymous):

Plus or minus.

OpenStudy (anonymous):

So it's the third one.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

You're welcome!